# On Modeling and Solving the Boltzmann Equation

### v1

@@ 6 The Integrodifferential Formulation: Neutrons
where now $\Psi({\bf r},{\bm{\Omega}})$ is the angular flux at ${\bf r}=(x,y,z)$ along direction $\bm{\Omega}$}; ${\bm{\Omega}}=(\mu,\eta,\xi)$ represents the direction of the particle as a vector on the unit sphere $S$; $\sigma_{t}$ represents the total macroscopic cross section; $\sigma_{s}({\bf r},{\bm{\Omega}^{\prime}}\cdot{\bm{\Omega}})$ is referred to as the differential scattering macroscopic cross section; $Q({\bf r},{\bm{\Omega}})$ is the fixed neutron source term. As for the RTE case discussed in Sec. (2), the equation will be suplemmented with boundary conditions.
In fact, the first works on the application of the ADO method to two-dimensional geometry problems were devoted to neutron transport (Barichello et al., 2011; Picoloto et al., 2015). Much work has been done with respect to the solution of a known benchmark problem in reactor shielding applications. We describe the geometry of the problem in Fig. (7). Comparisons (Barichello et al., 2017) with alternative nodal schemes available in the literature have shown a better perfomance of the ADO method in coarse meshes, as well as, expressive gain in computational time.
@@
However, the idea of bringing these applications to the attention here, has to do with the analysis of assymptotic error of the spatial discretization of the discrete-ordinates approximation of the two-dimensional neutron transport equation.
Madsen (1971) discussed the pointwise convergence of the three-dimensional discrete ordinates method, showing, in particular, the relation and the relevance of the truncation error in the scattering kernel approximation. We worked on a detailed analysis on that (Tres, 2015; Barichello et al., 2016; Hu et al., 2016; Nenê, 2020).
Using Richardson Extrapolation (Richardson and Gaunt, 1927), a spatial asymptotic analysis was made (Barichello et al., 2016) in order to determine the convergence order of nodal methods and a spatial reference solution. Numerical results for region-averaged scalar flux were generated for comparisons between the ADO method and the AHOT-N0 (Azmy, 1992) another nodal technique.
We investigated a model problem which consists of a homogeneous square domain with sides a=b=1cma=b=1~cm, with a source $Q(x,y)=1.0$ in the region $0,0.5]\times[0,0.5]$, as shown in Fig. (7). The total cross section in the homogeneous domain is $\sigma_{t}=1.0~~cm^{-1}$ and we consider two different configurations where the scattering cross section assumes two values: $\sigma_{s}=0.9~~cm^{-1}$ in the first case and $\sigma_{s}=0.3~~cm^{-1}$ in the other.
We consider each one of the four regions as a node, when implementing the ADO method, and, to analyze the spatial convergence aspect we utilize $2^{j}\times 2^{j}$, $j=1,\ldots,8$, uniform mesh, when using the AHOT-N0 method. We investigate the spatial-discretization convergence through the application of the four quadrature schemes: $LQ_{N}$ (Lewis and Miller, 1984), $P_{N}T_{N}S_{N}$ (Longoni and Haghighat, 2001), $P_{N}T_{N}$ (Cacuci, 2010) and QR (Abu-Shumays, 1977) to the sequence of refined meshes in AHOT-N0. The maximum number of directions per octant (ND) considered was $105$. Details are given in Barichello et al. (2016) including tables listing the numerical results. Here we reproduce some results given in that work, in Figs. (8), (9), and (10).
We proceeded (Barichello et al., 2016) with a spatial asymptotic analysis in order to determine the convergence order $p$ and a spatial reference solution. For this purpose, we use the AHOT-N0 (Azmy, 1992) results and Richardson extrapolation (Richardson and Gaunt, 1927) to estimate an asymptotically extrapolated reference solution. For a given number of directions per octant, ND, we assume $\phi(h)$ as a numerical approximation of order $p$ to the extrapolated (reference) solution, $\phi_{ref}$(ND). In this way, we write
@@
and take evaluations on $h$, rhrh and $r^{2}h$, , where $r<1$ is a fixed factor,
@@
$\phi_{ref}=\left(\dfrac{r^{p}\phi(rh)-\phi(h}{r^{p}-1}\right).$ (97)
Figure 8: Relative Error Computed Against Spatially-extrapolated Reference Solution: Quadratic Convergence with Mesh Refinement for AHOT: points for shown quadratures lie atop one another (Region 1)
Figure 9: Scalar flux absolute error against reference solution, Regions2 and 3.
Figure 10: Scalar flux absolute error against reference solution, Regions2 and 4.
Results shown as $p$ order and resulting reference solution were calculated from the three finest meshes data with all digits available. We found that the absolute error computed against the spatially-extrapolated reference solution for AHOT-N0 results, for the source region, exhibits two features: (i) quadratic convergence with mesh refinement, (ii) that is practically independent of the quadrature choice. These two features are illustrated in Fig. (8).
On the other hand, for Regions $2,3$ and $4$, a tighter convergence criterion of order $10^{-11}$ for refined meshes ($128\times 128$ and $256\times 256$) were required to indicate quadratic convergence order. Still, the behavior for each quadrature is somewhat different (see Figs. (9) and (10)). We found from the results that asymptotic convergence in a strict sense ($p$ approaching a constant value) is not clearly established. The reason for this behavior is not evident and has deserved further investigation, nevertheless the fact that for the vast majority of cases considered the computed $p\sim 2$ is satisfactory.
@@ 6.1 Numerical Results and Contributions
The ADO formulation, extended to the solution of two-dimensional problems through the use of nodal schemes, is referred to as ADO-Nodal and is distinguished from other nodal schemes in that it retains explicit expressions for the average intensities in terms of each spatial variable. In addition, and very significantly, it has been associated with a class of quadrature schemes (another notable feature), retaining the order of the eigenvalue problem at half the number of directions. In the derivation presented in Section 2, the eigenvalue problem was extended to allow the formulation to be applied to problems with an arbitrary degree of anisotropy using both the expanded and exact formulation. It represents another feature, as far as we know, singular in comparison with nodal schemes available in the literature – without forgetting the issue that scattering is a significant challenge in these models. Numerical results that confirm such analysis may be found in the works already listed in Sections 2 to 6 and to appear.
@@ 7 A Taste of the Linearized Boltzmann Equation
Internal rarefied gas flows define a field of major interest in the general area of rarefied gas dynamics (Cercignani, 1969, 1988). We have seen an increased interest in this are because of applications, for example, as related to micro-eletro-mechanical machines (MEMS), where the Boltzmann equation or a model equation is required in order to describe well gas-flow and heat-flow mechanisms. In many cases the flow conditions are in the transition regime and as a result the well known and commonly used Navier-Stokes equations can not be applied. In these cases the Boltzmann equation or suitable kinetic models should be utilized.
The ADO method (Barichello and Siewert, 1999b) has been shown (Barichello et al., 2001; Barichello and Siewert, 2000b; Barichello et al., 2002a) to be a useful method for solving a class of basic problems in this field. For example, many of the classical problems based on the BGK model (Bhatnagar et al., 1954) have been solved in a unified and especially accurate way with the ADO method. Following beginning work with the BGK model, we extended the ADO method in order to solve many of the basic problems in rarefied gas dynamics that were defined in terms of the variable collision frequency model (CLF model) of Cercignani (1966) and (Loyalka and Ferziger, 1967). Having seen that the ADO method was a convenient computational method for solving problems based on the CLF model and seeking improved results for physical quantities of interest, we moved forward (Barichello and Siewert, 2003) and defined explicitly two extended model equations (CES model and the CEBS model) based on approximated forms of the LBE. We made use of exact solutions of the homogeneous and of the inhomogeneous LBE to provide a systematic basis for approximating the scattering kernel in the LBE.
Although intensive numerical methods can be used, as can the Monte Carlo method, to solve practical problems based on the LBE, for rigid-sphere collisions, our goal was to explore a class of model equations that can be solved (essentially) analytically to yield results good enough for selected applications and for producing bechmark solutions. We provided a systematic derivation of kinetc models and report the final forms in sufficient detail that researchers seeking to develop numerical algorithms to solve practical problems could have a clearly defined starting point.
In Barichello and Siewert (2003) we follow the papers of Pekeris and Alterman (1957) and Loyalka and Hickey (1989) and consider the LBE written as
@@
In regard to Eq. (98), we state that (Pekeris and Alterman, 1957; Loyalka and Hickey, 1989) $\sigma_{0}$ is the collision diameter of the gas particles (in the rigid-sphere approximation) and that the collision process is described by
${L}\{h\}(x,\bm{c})=-\nu(c)h(x,\bm{c})+\int_{0}^{\infty}\int_{-1}^{1}\int_{0}^{2\pi}{\text{e}}^{-{c^{\prime}}^{2}}h(x,\bm{c}^{\prime})K(\bm{c}^{\prime},\bm{c}){c^{\prime}}^{2}\,\text{d}\chi^{\prime}\,\text{d}\mu^{\prime}\,\text{d}c^{\prime},$ (101)
where we use spherical coordinates ($c,\arccos\mu,\chi$) to define the (dimensionless) velocity vector, where $\nu(c)$ is the “collision frequency” and where $K(\bm{c}^{\prime},\bm{c})$ is the scattering kernel. It is clear that to proceed we require definitions of the scattering kernel $K(\bm{c}^{\prime},\bm{c})$ and the collision frequency $\nu(c)$. In their work, Pekeris and Alterman (1957) used an expansion in terms of Legendre polynomials (as functions of the scattering angle between “before” and “after” directions) to describe the scattering process. We use the spherical-harmonics addition theorem and write the scattering kernel as
@@
$P_{n}^{m}(\mu)=\left[\frac{(n-m)\,!}{(n+m)\,!}\right]^{1/2}{(1-\mu^{2})}^{m/2}\frac{\,\text{d}^{m}\,\,}{\,\text{d}\mu^{m}}P_{n}(\mu),\quad n\geq m,$ (103)
@@
$\int_{-1}^{1}P_{n}^{m}(\mu)P_{n^{\prime}}^{m}(\mu)\,\text{d}\mu=\Big{(}\frac{2}{2n+1}\Big{)}\delta_{n,n^{\prime}}.$ (104)
The basic elements $k_{n}(c^{\prime},c)$ in the expansion of the scattering kernel are available from the paper of Pekeris and Alterman (1957) where explicit expressions are given for $n=$ 1 and 2.
@@
To have our final results in terms of the real spatial variable $x$ (in $cm$), then what we use for a mean-free path $l$ is not important if the molecular diameter $\sigma_{0}$ and the density $n_{0}$ are known. However, since these physical quantities may not be known some researchers choose to work in terms of a specific mean-free path. We can mention two convenient choices. For the Poiseuille-flow problem, we follow Loyalka and Hickey (1989) and found that (Barichello and Siewert, 2003),
$\varepsilon=\varepsilon_{\text{p}}=\frac{16}{15}\pi^{-1/2}\int_{0}^{\infty}{\text{e}}^{-c^{2}}b(c)c^{6}\,\text{d}c.$ (107)
For the temperature-jump problem, following Loyalka and Ferziger (1967) we use
$\varepsilon=\varepsilon_{\text{t}}=\frac{16}{15}\pi^{-1/2}\int_{0}^{\infty}{\text{e}}^{-c^{2}}a(c)c^{6}\,\text{d}c.$ (108)
In regard to numerical work, we note that Hermite cubic splines have been used (Barichello et al., 2004) to solve the Chapman-Enskog integral equations for viscosity and heat conduction, making possible to evaluate
$\varepsilon_{\text{p}}=0.449027806...\quad{\text{and}}\quad\varepsilon_{\text{t}}=0.679630049.$..\,. (109)
Noting that the Prandtl number normally used in kinetic theory (in contrast to the differing by 5/2 definition used in fluid dynamics) written as
${\text{Pr}}=\frac{5}{2}\frac{\mu_{*}k}{m\lambda_{*}}$ (110)
@@
${\text{Pr}}=\varepsilon_{\text{p}}/\varepsilon_{\text{t}},$ (111)
and so using Eqs. (109) in Eq. (111), we find the result
${\text{Pr}}=0.660694457$... (112)
@@ 7.1 A Synthetic Kernel
Even if we truncate the expansion of the scattering kernel given as Eq. (102) after only a few terms, the problem of solving the resulting approximation of the LBE is still difficult from a numerical point-of-view. The numerical difficulty comes about basically because the components $k_{n}(c^{\prime},c)$ required in Eq. (102) have derivatives (even for small values of $n$) that are discontinuous at $c^{\prime}=c.$ Keeping in mind that we intend to implement our work numerically, we seek to approximate the true kernel by physically meaningful approximations that can be more easily incorporated into a numerical algorithm. The variable collision frequency (CLF) model, for instance, (Cercignani, 1966) and (Loyalka and Ferziger, 1967), was used in Barichello and Siewert (2000b); Barichello et al. (2002a) in order to try to improve basic results available from the classical BGK model (Bhatnagar et al., 1954). In (Barichello and Siewert, 2003) we extend the variant of the variable collision frequency model used in Refs. (Barichello and Siewert, 2000b; Barichello et al., 2002a) by replacing the exact components $k_{n}(c^{\prime},c)$ by a truncated form of the scattering kernel $K(\bm{c}^{\prime},\bm{c})$ with a more general synthetic approximation. To begin, we truncate Eq. (102) and write the synthetic scattering kernel as
@@
${L^{*}}\{h\}(\tau,\bm{c})=-\nu(c)h(\tau,\bm{c})+\int_{0}^{\infty}\int_{-1}^{1}\int_{0}^{2\pi}{\text{e}}^{-{c^{\prime}}^{2}}h(\tau,\bm{c}^{\prime})F(\bm{c}^{\prime},\bm{c}){c^{\prime}}^{2}\,\text{d}\chi^{\prime}\,\text{d}\mu^{\prime}\,\text{d}c^{\prime}.$ (118)
@@
The definition of the components $f_{n}(c^{\prime},c),n=0,1,2,$ of the $F(\bm{c}^{\prime},\bm{c})$, detailed in Barichello and Siewert (2003) results from the requirement that exact solutions of the LBE satisfy also the model equation. This derivation involves the solution of a class of integral equations (Barichello et al., 2004). Such formulation allowed us to propose two additional kinetic models, referred to as CES and CEBS, to the ones available in the literature by that time (BGK, S-model, CLF).
In defining what we have called the CES model of the linearized Boltzmann equation, we established five conditions on the synthetic kernel $F(\bm{c}^{\prime},\bm{c})$ by insisting that the known solutions of the homogeneous LBE be also solutions of the model equation. As we have no more known solutions we require other conditions to have a more general model equation. Since the classical problems of Poiseuille flow and thermal-creep flow in a plane channel are normally defined in terms of an inhomogeneous version of the LBE, we define our new model equation by insisting that the particular solution required for the LBE be also a particular solution of the inhomogeneous model equation. With that, we determined $f_{n}(c^{\prime},c),n=0,1,2,3$, so as to define a new synthetic kernel. The CEBS model has one more term ($N=3$ rather than $N=2$) in the synthetic scattering kernel, and it proved to be a good addition to the class of model equations already available for approximating the linearized Boltzmann equation.
In summary, as we state in Barichello and Siewert (2003), in our work on the rarefied gas dynamics, we have made consistent use of the idea of approximating the exact scattering kernel relevant to the linearized Boltzmann equation for rigid-sphere interactions with a synthetic kernel that maintains some basic properties of the exact kernel. To be clear, we insist that a model equation defined by an approximating synthetic kernel accept as solutions certain known solutions of the homogeneous LBE or the inhomogeneous LBE relevant to forced flow in a plane channel. By developing improved model equations (what we call the CES model and the CEBS model) that are amenable to analytical and simple numerical methods of solution, we provided conditions that numerical results for applications of interest can be made available without the need of extensive computation methods required when the LBE must be solved numerically. In addition to that, it allowed us to model and solve in an unified way a series of classical problems in RGD: using the same methodology, comparing the results of different kinetic models providing concise and accurate solutions for such problems, including gas mixtures and the analysis of the Cercignani-Lampis boundary condition. Some of these results can be found in Cabrera and Barichello (2006); Knackfuss and Barichello (2006); Scherer et al. (2009); Scherer and Barichello (2010); Siewert (2004b).

### v2

@@ 6 Neutron Transport
where now $\Psi({\bf r},{\bm{\Omega}})$ is the angular flux at ${\bf r}=(x,y,z)$ along direction $\bm{\Omega}$}; ${\bm{\Omega}}=(\mu,\eta,\xi)$ represents the direction of the particle as a vector on the unit sphere $S$; $\sigma_{t}$ represents the total macroscopic cross section; $\sigma_{s}({\bf r},{\bm{\Omega}^{\prime}}\cdot{\bm{\Omega}})$ is referred to as the differential scattering macroscopic cross section; $Q({\bf r},{\bm{\Omega}})$ is the fixed neutron source term. As for the RTE case discussed in Sec. (2), the equation will be supplemented with boundary conditions.
In fact, the first works on the application of the ADO method to two-dimensional geometry problems were devoted to neutron transport [6, 53]. Much work has been done with respect to the solution of a known benchmark problem in reactor shielding applications. We describe the geometry of the problem in Fig. 7. Comparisons [12] with alternative nodal schemes available in the literature have shown a better performance of the ADO method in coarse meshes, as well as an expressive gain in computational time.
@@
However, the idea of bringing these applications to the attention here has to do with the analysis of the asymptotic error of the spatial discretization of the discrete-ordinates approximation of the two-dimensional neutron transport equation.
Madsen [47] discussed the pointwise convergence of the three-dimensional discrete ordinates method, showing, in particular, its relation with the truncation error in the scattering kernel approximation. We worked on this direction [64, 21, 41, 49].
Using Richardson Extrapolation [55], a spatial asymptotic analysis was made [21] in order to determine the convergence order of nodal methods and a spatial reference solution. Numerical results for region-averaged scalar flux were generated for comparisons between the ADO method and the AHOT-N0 [2], another nodal technique.
We investigated a model problem which consists of a homogeneous square domain with sides a=b=1cma=b=1~cm, with a source $Q(x,y)=1.0$ in the region $0,0.5]\times[0,0.5]$, as shown in Fig. 7. The total cross section in the homogeneous domain is $\sigma_{t}=1.0~~cm^{-1}$ and we consider two different configurations where the scattering cross section assumes two values: $\sigma_{s}=0.9~~cm^{-1}$ in the first case and $\sigma_{s}=0.3~~cm^{-1}$ in the other.
We consider each one of the four regions as a node when implementing the ADO method, and to analyze the spatial convergence aspect we utilize $2^{j}\times 2^{j}$, $j=1,\ldots,8$, uniform mesh, when using the AHOT-N0 method. We investigate the spatial-discretization convergence through the application of the four quadrature schemes: $LQ_{N}$ [43], $P_{N}T_{N}S_{N}$ [44], $P_{N}T_{N}$ [26] and QR [1] to the sequence of refined meshes in AHOT-N0. The maximum number of directions per octant (ND) considered was $105$. Details are given in [21], including tables listing the numerical results. Here we reproduce some results given in that work, in Figs. 8, 9, and 10.
We proceeded [21] with a spatial asymptotic analysis in order to determine the convergence order $p$ and a spatial reference solution. For this purpose, we use the AHOT-N0 [2] results and Richardson extrapolation [55] to estimate an asymptotically extrapolated reference solution. For a given number of directions per octant, ND, we assume $\phi(h)$ as a numerical approximation of order $p$ to the extrapolated (reference) solution, $\phi_{ref}$(ND). In this way, we write
@@
and take evaluations on $h$, rhrh and $r^{2}h$, where $r<1$ is a fixed factor,
@@
$\phi_{ref}=\left(\dfrac{r^{p}\phi(rh)-\phi(h)}{r^{p}-1}\right).$ (97)
Results shown as $p$ order, and the resulting reference solution were calculated from the three finest meshes of data with all digits available. We found that the absolute error computed against the spatially-extrapolated reference solution for AHOT-N0 results, for the source region, it exhibits two features: (i) quadratic convergence with mesh refinement, (ii) that is practically independent of the quadrature choice. These two features are illustrated, for $\sigma_{s}=0.9~~cm^{-1}$, in Fig. 8.
Figure 8: Relative Error Computed Against Spatially-extrapolated Reference Solution (Region 1)
On the other hand, for Regions $2,3$ and $4$, a tighter convergence criterion of order $10^{-11}$ for refined meshes ($128\times 128$ and $256\times 256$) were required to indicate quadratic convergence order. Still, the behavior for each quadrature is somewhat different (see Figs. 9 and 10).
Figure 9: Scalar flux absolute error against reference solution (Regions 2 and 3)
Figure 10: Scalar flux absolute error against reference solution (Region 4)
We found from the results that asymptotic convergence in a strict sense ($p$ approaching a constant value) is not clearly established. The reason for this behavior is not evident and has deserved further investigation, nevertheless the fact that for the vast majority of cases considered the computed $p\sim 2$ is satisfactory.
@@ 6.1 Numerical Results and Contributions
The ADO formulation, extended to the solution of two-dimensional problems through the use of nodal schemes, is referred to as ADO-Nodal and is distinguished from other nodal schemes in that it retains explicit expressions for the average intensities in terms of each spatial variable. In addition, and very significantly, it has been associated with a class of quadrature schemes (another notable feature), retaining the order of the eigenvalue problem at half the number of directions. In the derivation presented in Section 5.2, the eigenvalue problem was extended to allow the formulation to be applied to problems with an arbitrary degree of anisotropy using both the expanded and exact formulation. It represents another feature, as far as we know, singular in comparison with nodal schemes available in the literature – without forgetting the issue that scattering is a significant challenge in these models. Numerical results that confirm such analysis may be found in the works listed in this text.
@@ 7 A Taste of the Linearized Boltzmann Equation
Internal rarefied gas flows define a field of major interest in the general area of rarefied gas dynamics [30, 31]. We have seen an increased interest in this area because of applications, for example, as related to micro-electro-mechanical machines (MEMS), where the Boltzmann equation or a model equation is required in order to describe well gas-flow and heat-flow mechanisms. In many cases, the flow conditions are in the transition regime, and as a result, the well-known and commonly used Navier-Stokes equations can not be applied. In these cases, the Boltzmann equation or suitable kinetic models should be utilized.
The ADO method [17] has been shown [7, 19, 5] to be a useful method for solving a class of basic problems in this field. For example, many of the classical problems based on the BGK model [23] have been solved in a unified and especially accurate way with the ADO method. Following the beginning work with the BGK model, we extended the ADO method in order to solve many of the basic problems in rarefied gas dynamics that were defined in terms of the variable collision frequency model (CLF model) of Cercignani [29] and Loyalka and Ferziger [45]. Having seen that the ADO method was a convenient computational method for solving problems based on the CLF model and seeking improved results for physical quantities of interest, we moved forward [22] and defined explicitly two extended model equations (CES model and the CEBS model) based on approximated forms of the LBE. We made use of exact solutions of the homogeneous and of the inhomogeneous LBE to provide a systematic basis for approximating the scattering kernel in the LBE.
Although intensive numerical methods can be used, as can the Monte Carlo method, to solve practical problems based on the LBE, for rigid-sphere collisions, our goal was to explore a class of model equations that can be solved (essentially) analytically to yield results good enough for selected applications and for producing benchmark solutions. We provided a systematic derivation of kinetic models and report the final forms in sufficient detail that researchers seeking to develop numerical algorithms to solve practical problems could have a clearly defined starting point.
In [22], following [52, 46], we consider the LBE written as
@@
In regard to Eq. (98), we state that [52, 46] $\sigma_{0}$ is the collision diameter of the gas particles (in the rigid-sphere approximation) and that the collision process is described by
${L}\{h\}(x,\bm{c})=-\nu(c)h(x,\bm{c})+\int_{0}^{\infty}\int_{-1}^{1}\int_{0}^{2\pi}\operatorname{e}^{-{c^{\prime}}^{2}}h(x,\bm{c}^{\prime})K(\bm{c}^{\prime},\bm{c}){c^{\prime}}^{2}\operatorname{d}\chi^{\prime}\operatorname{d}\mu^{\prime}\operatorname{d}c^{\prime},$ (101)
where we use spherical coordinates ($c,\arccos\mu,\chi$) to define the (dimensionless) velocity vector, where $\nu(c)$ is the “collision frequency” and where $K(\bm{c}^{\prime},\bm{c})$ is the scattering kernel. It is clear that to proceed we require definitions of the scattering kernel $K(\bm{c}^{\prime},\bm{c})$ and the collision frequency $\nu(c)$. In their work, Pekeris and Alterman [52] used an expansion in terms of Legendre polynomials (as functions of the scattering angle between “before” and “after” directions) to describe the scattering process. We use the spherical-harmonics addition theorem and write the scattering kernel as
@@
$P_{n}^{m}(\mu)=\left[\frac{(n-m)\,!}{(n+m)\,!}\right]^{1/2}{(1-\mu^{2})}^{m/2}\frac{\operatorname{d}^{m}\,\,}{\operatorname{d}\mu^{m}}P_{n}(\mu),\quad n\geq m,$ (103)
@@
$\int_{-1}^{1}P_{n}^{m}(\mu)P_{n^{\prime}}^{m}(\mu)\operatorname{d}\mu=\Big(\frac{2}{2n+1}\Big)\delta_{n,n^{\prime}}.$ (104)
The basic elements $k_{n}(c^{\prime},c)$ in the expansion of the scattering kernel are available from the paper of [52] where explicit expressions are given for $n=$ 1 and 2.
@@
To have our final results in terms of the real spatial variable $x$ (in $cm$), then what we use for a mean-free path $l$ is not important if the molecular diameter $\sigma_{0}$ and the density $n_{0}$ are known. However, since these physical quantities may not be known, some researchers choose to work in terms of a specific mean-free path. We can mention two convenient choices. For the Poiseuille-flow problem, we follow [46] and found that [22],
$\varepsilon=\varepsilon_{\operatorname{p}}=\frac{16}{15}\pi^{-1/2}\int_{0}^{\infty}\operatorname{e}^{-c^{2}}b(c)c^{6}\operatorname{d}c.$ (107)
For the temperature-jump problem, following [45] we use
$\varepsilon=\varepsilon_{\operatorname{t}}=\frac{16}{15}\pi^{-1/2}\int_{0}^{\infty}\operatorname{e}^{-c^{2}}a(c)c^{6}\operatorname{d}c.$ (108)
In regard to numerical work, we note that Hermite cubic splines have been used [14] to solve the Chapman-Enskog integral equations for viscosity and heat conduction, making it possible to evaluate
$\varepsilon_{\operatorname{p}}=0.449027806...\quad{\text{and}}\quad\varepsilon_{\operatorname{t}}=0.679630049.$..\,. (109)
Noting that the Prandtl number normally used in kinetic theory (in contrast to the difference by 5/2 definition used in fluid dynamics) written as
$\Pr=\frac{5}{2}\frac{\mu_{*}k}{m\lambda_{*}}$ (110)
@@
$\Pr=\varepsilon_{\operatorname{p}}/\varepsilon_{\operatorname{t}},$ (111)
and so using Eq. (109) in Eq. (111), we find the result
$\Pr=0.660694457$... (112)
@@ 7.1 A Synthetic Kernel
Even if we truncate the expansion of the scattering kernel given as Eq. (102) after only a few terms, the problem of solving the resulting approximation of the LBE is still difficult from a numerical point of view. The numerical difficulty comes about basically because the components $k_{n}(c^{\prime},c)$ required in Eq. (102) have derivatives (even for small values of $n$) that are discontinuous at $c^{\prime}=c.$ Keeping in mind that we intend to implement our work numerically, we seek to approximate the true kernel by physically meaningful approximations that can be more easily incorporated into a numerical algorithm. The variable collision frequency (CLF) model, for instance, [29] and [45], was used in [19, 5] in order to try to improve basic results available from the classical BGK model [23]. In [22], we extend the variant of the variable collision frequency model used in [19, 5] by replacing the exact components $k_{n}(c^{\prime},c)$ by a truncated form of the scattering kernel $K(\bm{c}^{\prime},\bm{c})$ with a more general synthetic approximation. To begin, we truncate Eq. (102) and write the synthetic scattering kernel as
@@
${L^{*}}\{h\}(\tau,\bm{c})=-\nu(c)h(\tau,\bm{c})+\int_{0}^{\infty}\int_{-1}^{1}\int_{0}^{2\pi}\operatorname{e}^{-{c^{\prime}}^{2}}h(\tau,\bm{c}^{\prime})F(\bm{c}^{\prime},\bm{c}){c^{\prime}}^{2}\operatorname{d}\chi^{\prime}\operatorname{d}\mu^{\prime}\operatorname{d}c^{\prime}.$ (118)
@@
The definition of the components $f_{n}(c^{\prime},c),n=0,1,2,$ of the $F(\bm{c}^{\prime},\bm{c})$, detailed in [22], results from the requirement that exact solutions of the LBE satisfy also the model equation. This derivation involves the solution of a class of integral equations [14]. Such formulation allowed us to propose two additional kinetic models, referred to as CES and CEBS, to the ones available in the literature by that time (BGK, S-model, CLF).
In defining what we have called the CES model of the linearized Boltzmann equation, we established five conditions on the synthetic kernel $F(\bm{c}^{\prime},\bm{c})$ by insisting that the known solutions of the homogeneous LBE be also solutions of the model equation. As we have no more known solutions, we require other conditions to have a more general model equation. Since the classical problems of Poiseuille flow and thermal-creep flow in a plane channel are normally defined in terms of an inhomogeneous version of the LBE. We define our new model equation by insisting that the particular solution required for the LBE be also a particular solution of the inhomogeneous model equation. With that, we determined $f_{n}(c^{\prime},c),n=0,1,2,3$, so as to define a new synthetic kernel. The CEBS model has one more term ($N=3$ rather than $N=2$) in the synthetic scattering kernel, and it proved to be a good addition to the class of model equations already available for approximating the linearized Boltzmann equation.
In summary, as we state in [22], in our work on the rarefied gas dynamics, we have made consistent use of the idea of approximating the exact scattering kernel relevant to the linearized Boltzmann equation for rigid-sphere interactions with a synthetic kernel that maintains some basic properties of the exact kernel. To be clear, we insist that a model equation defined by an approximating synthetic kernel accept as solutions certain known solutions of the homogeneous LBE or the inhomogeneous LBE relevant to forced flow in a plane channel. By developing improved model equations (what we call the CES model and the CEBS model) that are amenable to analytical and simple numerical methods of solution, we provided conditions that numerical results for applications of interest can be made available without the need of extensive computation methods required when the LBE must be solved numerically. In addition to that, it allowed us to model and solve in a unified way a series of classical problems in RGD: using the same methodology, comparing the results of different kinetic models, providing concise and accurate solutions for such problems, including gas mixtures and the analysis of the Cercignani-Lampis boundary condition. Some of these results can be found in [25, 42, 58, 59, 61].